# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 2.2.2 Conversational Engagements in Healthcare
Healthcare is one of the key domains where multi-turn conversational large language models (LLMs) demonstrate significant potential. A defining characteristic of these models in the medical domain is their ability to emulate a doctor’s role by engaging in task-oriented, context-aware dialogues with patients. Unlike single-turn medical knowledge question-answering systems, such as BenTsao  [102] (formerly named HuaTuo), ChatMed, ShenNong-TCM, MING [103], and DoctorGLM [104], responding to isolated queries with full information at a time [105], multi-turn healthcare LLMs operate in a setting where patient information is often incomplete or ambiguous at the outset.
Ideally, multi-turn healthcare conversational LLMs should be capable of proactively generating a sequence of questions to refine their understanding of the patient’s condition through iterative inquiry. This concept has been referred to by different names in various studies, including chain of questions (CoQ) [106], proactive questioning [106], symptom inquiry [107, 108], proactivity [109], and information seeking [110]. Despite the differences in terminology, they all describe the model’s ability to dynamically gather relevant details over multiple exchanges, ultimately leading to a more accurate diagnosis. Additionally, multi-turn healthcare LLMs need to retain dialogue history, so that enabling seamless continuity in conversations, and integrate medical knowledge to ensure responses are accurate, reliable, and contextually appropriate.
While prior surveys [15, 16] provide a broad overview of medical dialogue systems, including both pre-LLM and LLM-based approaches, our focus is specifically on multi-turn LLMs and related evaluation framework. Traditional medical dialogue systems typically rely on rule-based logic or task-specific neural networks, which can be rigid and require extensive manual engineering. We explore the specific tasks these models address, the architectural advancements in their development, and the characteristics of the datasets used for pre-training and fine-tuning. In addition, we discuss the evaluation frameworks used to assess these models, highlighting key metrics, established benchmarks, the growing focus on their information-seeking capabilities, and the challenges faced in online conversational healthcare systems.
@@
A collection of studies have aimed to make medical LLM communication more interactive and accessible. For example, Clinical Camel [111] is an open-source medical language model that introduces dialogue-based knowledge encoding (DBKE) to transform dense medical texts into conversational formats. This methodology enhances the model’s ability to engage in multi-turn dialogues, aligning its responses to a conversational format. Fine-tuned from LLaMA-2 using QLoRA, Clinical Camel outperforms other LLMs, including GPT 3.5, GPT 4, and Med-PaLM2, across benchmarks such as USMLE Sample Exam [112], PubMedQA [113], and MedQA [114]. T-Agent [115] enhances medical dialogue generation by incorporating a term-aware approach by integrating a term extraction tool and a term prediction model within a two-stage training framework. This design aims to enhance the model’s performance in dialogue term status extraction and generation tasks. Their experimental result demonstrates the improvements of T-Agent in the ROUGE and term extraction F1 scores. In contrast, APP (Ask Patients with Patience) [116] serves as a reasoning and guidance layer atop an LLM to support diagnostic decision-making during online medical consultations. Rather than altering dialogue format, APP emphasizes grounded reasoning, entropy minimization, and patient-centered communication without requiring further fine-tuning.
@@
A majority of studies have focused on building medical LLMs that generate accurate and reliable responses in dialogue settings through SFT on healthcare-specific data. DISC-MedLLM [109] is fine-tuned on DISC-Med-SFT, a 400K-sample Chinese medical instruction dataset covering single-turn Q&A, multi-turn consultations, and multiple-choice Q&A, which is evaluated based on rule-based accuracy for single-turn and GPT-4 scoring for multi-turn conversations on Proactivity, Accuracy, Helpfulness, and Linguistic Quality. BianQue [106] integrates multi-turn doctor-patient Q&A datasets to enable a Chain of Questioning (CoQ) approach, emulating real consultations where doctors iteratively inquire to fully understand a patient’s condition. Fine-tuned on ChatGLM-6B using the 2.4M-sample BianQueCorpus, BianQue improves CoQ by balancing questions and suggestions, addressing earlier models’ limitations in interactive questioning. BiMediX [117], a bilingual medical mixture-of-experts LLM, enables seamless interaction in both English and Arabic. Fine-tuned using QLoRA, BiMediX outperformed existing models like Med42 and Meditron in English-based medical evaluations and significantly surpassed the generic bilingual LLM Jais-30B in Arabic medical and bilingual assessments. CPsyCounX [118] and PsycoLLM [119] are two healthcare conversational LLMs focusing on psychological counseling, where CPsyCounX is fine-tuned over InternLM2-7B-Chat with CPsyCounD, which contains 3,134 high-quality multi-turn consultation dialogues, and PsycoLLM is fine-tuned with psychological single-turn Q&A, multiturn dialogues, and knowledge-based Q&A on based on Qwen1.514B-Chat. Besides, the SMILE [120] method (Single-turn to Multi-turn Inclusive Language Expansion) uses ChatGPT to rewrite single-turn counseling QA exchanges into multi-turn dialogues. Starting from a mental health QA dataset (PsyQA [121]), SMILE injects diverse dialogue topics and prompts ChatGPT to produce realistic counselor–client conversations, then filters out any outputs that don’t meet format or turn-count requirements. This process yielded SMILECHAT, a corpus of 55k multi-turn counseling dialogues in Chinese
@@
Beyond solely relying on supervised fine-tuning (SFT), some studies have adopted a comprehensive pipeline encompassing pre-training, SFT, and reinforcement learning from human feedback (RLHF). Zhongjing [122], the first Chinese medical LLaMA-based large language model, exemplifies this approach through continuous pre-training, targeted SFT, and RLHF optimization. Particularly emphasizing multi-turn interactions, the authors constructed CMtMedQA, a specialized Chinese medical dialogue dataset comprising approximately 7,000 QA pairs derived from authentic doctor-patient exchanges across 14 clinical departments and spanning over 10 medical scenarios, including disease diagnosis, medication guidance, health consultation, and general medical knowledge inquiries. To address the inconsistency and brevity prevalent in real-world doctor responses, a self-instruct methodology [123] was utilized, standardizing replies into a uniform, professional, and empathetic communication style. Furthermore, the external medical knowledge graph CMeKG [124] was integrated to verify and enhance the medical accuracy and safety of the dialogue interactions. HuaTuoGPT [107] is a medical consultation LLM, fine-tuned using both data distilled from ChatGPT and real-world data from medical professionals to combine fluent and informative responses with authentic diagnostic capabilities. The authors employ Reinforcement Learning from AI Feedback to align the model’s outputs with the strengths of both data sources. HuatuoGPT is evaluated using rule-based NLP metrics for similarity check with reference answers, LLM-based assessment for language quality, symptom inquiry, treatment effectiveness, and patient helpfulness, and human evaluation by medical experts for diagnosis accuracy, treatment recommendation accuracy, and prescription knowledge, which demonstrates superior performance compared to baseline models under different criteria. The same research team later developed HuaTuoGPT II [125], which adopts a one-stage domain adaptation protocol that unifies heterogeneous data from traditional pre-training and supervised stages into a simple instruction-output pair format, facilitating efficient knowledge injection. The model achieved competitive performance with GPT-4 across multiple benchmarks, notably excelling in Chinese medical evaluations and the latest pharmacist license examinations.
@@
Several papers also apply Direct Preference Optimization (DPO) within the full pipeline to generate responses that better align with human preferences. Aquila-Med [126] is a medical LLM that undergoes continued pre-training, SFT, and reinforcement learning with DPO to enhance its medical consultation capabilities. 12,727 DPO preference pairs are utilized for fine-tuning, which demonstrates significant improvements in handling single-turn and multi-turn medical consultations, indicating enhanced fluency, relevance, completeness, and proficiency. Similarly, Qilin-Med [127] utilizes a multi-stage training approach that combines Domain-specific Continued Pre-training (DCPT), SFT, and DPO. This method leverages the ChiMed dataset [128], encompassing question answering, plain texts, knowledge graphs, and dialogues, resulting in an accuracy improvement with the baseline Baichuan-7B model on the CMExam [129] test set during the SFT phase. In the DPO phase, Qilin-Med achieved scores of 16.66 in BLEU-1 and 27.44 in ROUGE-1 on the Huatuo-26M test set, indicating further enhancements over the SFT phase.
@@
In addition to the mainstream approaches, Google’s Articulate Medical Intelligence Explorer (AMIE) [130] represents a comprehensive advancement in conversational medical AI for training and evaluation. AMIE utilizes a chain-of-reasoning strategy within a simulated environment, incorporating self-play and automated feedback mechanisms to enhance its diagnostic dialogue capabilities across diverse medical conditions. Its training data includes both human-curated and AI-generated datasets, encompassing multiple-choice medical questions, long-form medical reasoning queries, clinical note summaries, and simulated dialogues based on various medical conditions. Evaluated through a randomized, double-blind crossover study using Objective Structured Clinical Examination scenarios, AMIE demonstrated superior diagnostic accuracy compared to primary care physicians and received higher ratings from both specialist physicians and patient actors on multiple assessment criteria.
@@
Recent studies have proposed evaluation frameworks for medical LLMs that emphasize interactive quality, clinical reasoning, and human-centered communication. MedGPTEval [131] assesses LLMs like ChatGPT and Dr.PJ using 27 multi-turn dialogue cases and 7 case reports, measuring the accuracy, empathy, and clinical logic of the LLMs’ responses. Similarly, Liao et al. (2023) [132] proposes an automated evaluation framework, emphasizing assessing LLM abilities, such as recognizing knowledge limitations, gathering relevant information, and improving diagnostic accuracy. The researchers reformulated medical multiple-choice questions from the USMLE into consultation tasks, creating a specialized benchmark for assessment. Additionally, they verify that fine-tuning with a consultation-specific dataset reduced hallucinations and improved benchmark performance.
@@
Besides, a growing number of studies emphasize simulation-based interactive evaluation to approximate real-world clinical consultation. For instance, Liao et al. (2024) [108] introduced the Automated Interactive Evaluation (AIE) framework featuring the State-Aware Patient Simulator (SAPS), which incorporates a state tracker, memory bank, and response generator to support dynamic, multi-turn evaluation. Similarly, MMD-Eval (Multi-turn Medical Dialogue Evaluation) [133] offers a locally deployable, task-oriented dialogue simulator, providing more resource-efficient and consistent evaluations than LLM-based scoring. MediQ [110] introduces an interactive benchmark for medical evaluation by simulating clinical interactions between a patient system and an adaptive expert system, emphasizing the assessment of information seeking ability. It employs abstention strategies to better estimate confidence and determine when to seek additional information. The benchmark converts existing datasets like MedQA and Craft-MD into interactive formats, simulating clinical interactions for evaluation purposes.
@@
Complementing these efforts, MedFuzz [134] focuses on evaluating the adversarial robustness of healthcare LLMs by introducing ambiguous or unexpected inputs that could lead to clinical misconceptions and harmful decisions. It questions the assumption that high benchmark scores reflect real-world reliability by introducing complexities like ambiguous patient traits and biased data into medical QA benchmarks. Results show that GPT-3.5, GPT-4, and Med-PaLM 2 perform worse on the “MedFuzzed” benchmark, showing their vulnerability to clinical biases, demographic stereotypes, and incomplete data interpretation.
@@
Recently, OpenAI introduced HealthBench [135], an open-source benchmark constructed from real-world healthcare conversations designed explicitly for evaluating LLM performance. HealthBench includes 5,000 multi-turn conversations averaging 2.6 turns per dialogue (ranging from 1 to 19 turns), capturing diverse and realistic patient-clinician interactions. It emphasizes clinical accuracy, appropriate communication depth, handling of uncertainty, and context awareness. Unique to HealthBench is its extensive use of physician-developed rubrics, consisting of over 48,000 distinct evaluation criteria, making it a robust tool for assessing nuanced aspects of clinical dialogue. Moreover, HealthBench introduces specialized subsets such as HealthBench Consensus, validated across multiple experts, and HealthBench Hard, comprising notably challenging interactions that test the limits of current LLM capabilities. This systematic, clinician-validated approach significantly enriches the landscape of multi-turn medical LLM evaluation by explicitly aligning automated assessment closely with clinical judgment.
@@
In the development of healthcare conversational LLMs, a widely adopted strategy is to construct multi-turn medical dialogue datasets and fine-tune models using SFT, sometimes followed by RL-based methods. Interestingly, while some studies in general multi-turn dialogue suggest that SFT+RLHF may not yield significant improvements, this pipeline has proven surprisingly effective in the healthcare domain. This may be attributed to the relatively high quality and domain specificity of medical data, which contrasts with the noisier, less structured data used in broader multi-turn applications. Evaluation metrics in this area include traditional rule-based NLP measures (e.g., BLEU and ROUGE), diagnostic accuracy, and both AI-based and human assessments using more nuanced, subjective, domain-knowledge judgment criteria. However, very few papers conduct the human and AI agreement check [130, 108, 125], even if a considerable of literature use both of them. Notably, information-seeking and clinical reasoning are essential capabilities in this domain and are increasingly prioritized in newer benchmarks, which now place more emphasis on clinical relevance than conventional NLP metrics. Please refer to the appendix for details on the papers discussed.

### v5

@@ 2.2.2 Conversational Engagements in Healthcare
Healthcare is one of the key domains where multi-turn conversational large language models (LLMs) demonstrate significant potential. A defining characteristic of these models in the medical domain is their ability to emulate a doctor’s role by engaging in task-oriented, context-aware dialogues with patients. Unlike single-turn medical knowledge question-answering systems, such as BenTsao  Wang et al. (2023a) (formerly named HuaTuo), ChatMed, ShenNong-TCM, MING Liao et al. (2024a), and DoctorGLM Xiong et al. (2023), responding to isolated queries with full information at a time Mutabazi et al. (2021), multi-turn healthcare LLMs operate in a setting where patient information is often incomplete or ambiguous at the outset.
Ideally, multi-turn healthcare conversational LLMs should be capable of proactively generating a sequence of questions to refine their understanding of the patient’s condition through iterative inquiry. This concept has been referred to by different names in various studies, including chain of questions (CoQ) (Chen et al., 2023), proactive questioning (Chen et al., 2023), symptom inquiry (Zhang et al., 2023; Liao et al., 2024b), proactivity (Bao et al., 2023), and information seeking (Li et al., 2024). Despite the differences in terminology, they all describe the model’s ability to dynamically gather relevant details over multiple exchanges, ultimately leading to a more accurate diagnosis. Additionally, multi-turn healthcare LLMs need to retain dialogue history, so that enabling seamless continuity in conversations, and integrate medical knowledge to ensure responses are accurate, reliable, and contextually appropriate.
While prior surveys Valizadeh and Parde (2022); Shi et al. (2024c) provide a broad overview of medical dialogue systems, including both pre-LLM and LLM-based approaches, our focus is specifically on multi-turn LLMs and related evaluation framework. Traditional medical dialogue systems typically rely on rule-based logic or task-specific neural networks, which can be rigid and require extensive manual engineering. We explore the specific tasks these models address, the architectural advancements in their development, and the characteristics of the datasets used for pre-training and fine-tuning. In addition, we discuss the evaluation frameworks used to assess these models, highlighting key metrics, established benchmarks, the growing focus on their information-seeking capabilities, and the challenges faced in online conversational healthcare systems.
@@
Several studies have aimed to make medical LLM communication more interactive and accessible. Clinical Camel (Toma et al., 2023) is an open-source medical LLM that introduces dialogue-based knowledge encoding (DBKE) to transform dense medical texts into conversational formats, enhancing multi-turn dialogue capability. Fine-tuned from LLaMA-2 via QLoRA, Clinical Camel outperforms GPT-3.5, GPT-4, and Med-PaLM2 on benchmarks including USMLE (FSMB, 2023), PubMedQA (Jin et al., 2019), and MedQA (Jin et al., 2021). T-Agent (Hu et al., 2024) enhances medical dialogue generation through a term-aware approach, integrating a term extraction tool and a term prediction model within a two-stage training framework; experiments demonstrate improvements in ROUGE scores and term extraction F1. In contrast, APP (Ask Patients with Patience) (Zhu et al., 2025) operates as a reasoning and guidance layer atop an existing LLM to support diagnostic decision-making in online consultations, emphasizing grounded reasoning, entropy minimization, and patient-centered communication without requiring additional fine-tuning.
@@
A majority of studies have focused on building medical LLMs that generate accurate and reliable responses in dialogue settings through SFT on healthcare-specific data. DISC-MedLLM (Bao et al., 2023) is fine-tuned on DISC-Med-SFT, a 400K-sample Chinese medical instruction dataset covering single-turn Q&A, multi-turn consultations, and multiple-choice Q&A, evaluated via rule-based accuracy for single-turn and GPT-4 scoring for multi-turn conversations across proactivity, accuracy, helpfulness, and linguistic quality. BianQue (Chen et al., 2023) integrates multi-turn doctor–patient Q&A to enable a Chain of Questioning (CoQ) approach that emulates real consultations, fine-tuned on ChatGLM-6B using the 2.4M-sample BianQueCorpus, improving CoQ by balancing questions and suggestions. BiMediX (Pieri et al., 2024), a bilingual mixture-of-experts LLM supporting English and Arabic, is fine-tuned via QLoRA and outperforms Med42 and Meditron on English medical benchmarks while substantially surpassing the bilingual Jais-30B on Arabic and bilingual assessments. In psychological counseling, CPsyCounX (Zhang et al., 2024a) is fine-tuned over InternLM2-7B-Chat on CPsyCounD—3,134 high-quality multi-turn consultation dialogues—while PsycoLLM (Hu et al., 2025b) is fine-tuned on single-turn Q&A, multi-turn dialogues, and knowledge-based Q&A atop Qwen1.5-14B-Chat. Complementing these, the SMILE method (Qiu et al., 2024) leverages ChatGPT to convert single-turn counseling pairs from PsyQA (Sun et al., 2021) into multi-turn dialogues via diverse topic prompts and format filtering, yielding SMILECHAT, a corpus of 55K multi-turn counseling dialogues in Chinese.
@@
Beyond SFT alone, several studies adopt a comprehensive pipeline encompassing pre-training, SFT, and RLHF. Zhongjing (Yang et al., 2024), the first Chinese medical LLaMA-based LLM, exemplifies this through continuous pre-training, targeted SFT, and RLHF optimization. The authors construct CMtMedQA—approximately 7,000 QA pairs from authentic doctor–patient exchanges across 14 clinical departments and over 10 medical scenarios—and apply a self-instruct methodology (Wang et al., 2023b) to standardize responses into a uniform, professional, and empathetic style. The external medical knowledge graph CMeKG (Byambasuren et al., 2019) is further integrated to verify medical accuracy and safety. HuaTuoGPT (Zhang et al., 2023) fine-tunes a medical consultation LLM on data distilled from ChatGPT and real-world clinical data, using RLAIF to align outputs with the complementary strengths of both sources. Evaluation spans rule-based NLP metrics, LLM-based assessment of language quality and diagnostic utility, and expert human evaluation of diagnosis and treatment accuracy, demonstrating superior performance over baselines. The same team subsequently developed HuaTuoGPT II (Chen et al., 2024f), which adopts a one-stage domain adaptation protocol that unifies heterogeneous pre-training and supervised data into a unified instruction-output format for efficient knowledge injection, achieving performance competitive with GPT-4 on multiple benchmarks and excelling in Chinese medical and pharmacist license examinations.
@@
Several studies also incorporate Direct Preference Optimization (DPO) into the full training pipeline. Aquila-Med (Zhao et al., 2024) undergoes continued pre-training, SFT, and DPO using 12,727 preference pairs, demonstrating significant improvements in fluency, relevance, completeness, and proficiency across single- and multi-turn medical consultations. Similarly, Qilin-Med (Ye et al., 2023) employs a multi-stage pipeline combining domain-specific continued pre-training, SFT, and DPO over the ChiMed dataset (Tian et al., 2019)—encompassing question answering, plain text, knowledge graphs, and dialogues. SFT improves accuracy over the Baichuan-7B baseline on CMExam (Liu et al., 2023), while the subsequent DPO phase achieves BLEU-1 of 16.66 and ROUGE-1 of 27.44 on the Huatuo-26M test set, reflecting further gains beyond SFT.
@@
In addition to the mainstream approaches, Google’s Articulate Medical Intelligence Explorer (AMIE) (Tu et al., 2025) represents a comprehensive advancement in conversational medical AI for training and evaluation. AMIE utilizes a chain-of-reasoning strategy within a simulated environment, incorporating self-play and automated feedback mechanisms to enhance its diagnostic dialogue capabilities across diverse medical conditions. Its training data includes both human-curated and AI-generated datasets, encompassing multiple-choice medical questions, long-form medical reasoning queries, clinical note summaries, and simulated dialogues based on various medical conditions. Evaluated through a randomized, double-blind crossover study using Objective Structured Clinical Examination scenarios, AMIE demonstrated superior diagnostic accuracy compared to primary care physicians and received higher ratings from both specialist physicians and patient actors on multiple assessment criteria.
@@
Recent studies have proposed evaluation frameworks for medical LLMs that emphasize interactive quality, clinical reasoning, and human-centered communication. MedGPTEval (Xu et al., 2023) assesses LLMs like ChatGPT and Dr.PJ using 27 multi-turn dialogue cases and 7 case reports, measuring the accuracy, empathy, and clinical logic of the LLMs’ responses. Similarly, Liao et al. (2023) proposes an automated evaluation framework, emphasizing assessing LLM abilities, such as recognizing knowledge limitations, gathering relevant information, and improving diagnostic accuracy. The researchers reformulated medical multiple-choice questions from the USMLE into consultation tasks, creating a specialized benchmark for assessment. Additionally, they verify that fine-tuning with a consultation-specific dataset reduced hallucinations and improved benchmark performance.
@@
Besides, a growing number of studies emphasize simulation-based interactive evaluation to approximate real-world clinical consultation. For instance, Liao et al. (2024b) introduces the Automated Interactive Evaluation (AIE) framework featuring the State-Aware Patient Simulator (SAPS), which incorporates a state tracker, memory bank, and response generator to support dynamic, multi-turn evaluation. Similarly, MMD-Eval (Multi-turn Medical Dialogue Evaluation) (Liu et al., 2025c) offers a locally deployable, task-oriented dialogue simulator, providing more resource-efficient and consistent evaluations than LLM-based scoring. MediQ (Li et al., 2024) introduces an interactive benchmark for medical evaluation by simulating clinical interactions between a patient system and an adaptive expert system, emphasizing the assessment of information seeking ability. It employs abstention strategies to better estimate confidence and determine when to seek additional information. The benchmark converts existing datasets like MedQA and Craft-MD into interactive formats, simulating clinical interactions for evaluation purposes.
@@
Complementing these efforts, MedFuzz (Ness et al., 2024) focuses on evaluating the adversarial robustness of healthcare LLMs by introducing ambiguous or unexpected inputs that could lead to clinical misconceptions and harmful decisions. It questions the assumption that high benchmark scores reflect real-world reliability by introducing complexities like ambiguous patient traits and biased data into medical QA benchmarks. Results show that GPT-3.5, GPT-4, and Med-PaLM 2 perform worse on the “MedFuzzed” benchmark, showing their vulnerability to clinical biases, demographic stereotypes, and incomplete data interpretation.
@@
Recently, OpenAI introduced HealthBench (Arora et al., 2025), an open-source benchmark constructed from real-world healthcare conversations designed explicitly for evaluating LLM performance. HealthBench includes 5,000 multi-turn conversations averaging 2.6 turns per dialogue (ranging from 1 to 19 turns), capturing diverse and realistic patient-clinician interactions. It emphasizes clinical accuracy, appropriate communication depth, handling of uncertainty, and context awareness. Unique to HealthBench is its extensive use of physician-developed rubrics, consisting of over 48,000 distinct evaluation criteria, making it a robust tool for assessing nuanced aspects of clinical dialogue. Moreover, HealthBench introduces specialized subsets such as HealthBench Consensus, validated across multiple experts, and HealthBench Hard, comprising notably challenging interactions that test the limits of current LLM capabilities. This systematic, clinician-validated approach significantly enriches the landscape of multi-turn medical LLM evaluation by explicitly aligning automated assessment closely with clinical judgment.
@@
In the development of healthcare conversational LLMs, a widely adopted strategy is to construct multi-turn medical dialogue datasets and fine-tune models using SFT, sometimes followed by RL-based methods. Interestingly, while some studies in general multi-turn dialogue suggest that SFT+RLHF may not yield significant improvements, this pipeline has proven surprisingly effective in the healthcare domain. This may be attributed to the relatively high quality and domain specificity of medical data, which contrasts with the noisier, less structured data used in broader multi-turn applications. Evaluation metrics in this area include traditional rule-based NLP measures (e.g., BLEU and ROUGE), diagnostic accuracy, and both AI-based and human assessments using more nuanced, subjective, domain-knowledge judgment criteria. However, very few papers conduct the human and AI agreement check (Tu et al., 2025; Liao et al., 2024b; Chen et al., 2024f), even if a considerable of literature use both of them. Notably, information-seeking and clinical reasoning are essential capabilities in this domain and are increasingly prioritized in newer benchmarks, which now place more emphasis on clinical relevance than conventional NLP metrics. Please refer to the appendix for details on the papers discussed.